Bridgeport Industries changed from the double-declining-balance to the straight-line method in 2018 on all its equipment. There was no change in the assets’ salvage values or useful lives. Plant assets, acquired on January 2, 2015, had an original cost of $1,740,800, with a $99,200 salvage value and an 8-year estimated useful life. Income before depreciation expense was $275,200 in 2017 and $332,800 in 2018.

Partially correct answer. Your answer is partially correct. Try again.

Prepare the journal entry to record depreciation expense in 2018.

Starting with income before depreciation expense, prepare the remaining portion of the income statement for 2017 and 2018.
Answer #1

Double decline rate = 100/8*2 = 25%

Book value of 2017 = 1740800*75%*75%*75% = 734400

Depreciation expense for 2018 = (734400-99200/5) = 127040

Account title and explanation Debit Credit
Depreciation expense 127040
Accumulated depreciation 127040

Calculate following

2017 2018
Income before dep 275200 332800
Less: Dep -244800 -127040
Net income 30400 205760